OpenStudy (anonymous):

Medal!! ((math)) Which of the following is equivalent to five and four-ninths divided by seven? seven-ninths one and two-sevenths five-sevenths five and four sixty-thirds I'm thinking B ..

OpenStudy (anonymous):

\[5\frac{ 4 }{9 }=49/(9).......divide... by..7=..49/9*7..=49/63=7/9 \]